Find all distinct (real or complex) eigenvalues of A. Then find a basis for the eigenspace of A corresponding to each eigenvaurgent please,thanks
Answer #1

A= 8 -9 -6 -1 6 4 9 -15 -10/

(A - x0) is,

s -9 {-6 -1 9} 100\ 6 -15)-10 10 4 -10/ lo 01/

(8 -2 -9 -6 6 -1 -7 4 9 -15 -10 - 1

R2 + R2+ 3 - . R1

R3 6 R3+ 3 - - R1

18- -1 12-14r +39 8-1 -4r+26 8-1 9 151-39 8-1 12+2r-26 8- 1 0

R2 HR

x2 – 14.1 +39 R3 + R3-4 - 4.c + 26 - R

1 -4-26 8-C -1 9 | r2-27-26 | 0 上 8-I [00 - r(r=-4x+13) 2012-13) ST

now matrix is in triangular form

hence determinant is equal to the multiplication of diagonal entry

(8 - x) - 4.0 + 26 8 - (x2 - 4x + 13) 2 (2.– 13)

1 (-4.0 +26) (- 0:22 - 4x + 13) 2 2.– 13)

(-1 (22 – 4.2 + 13)) = 0

(22 – 4.1 +13) = 0

I=0, 12 – 4.0 + 13 = 0

2-4x+13=0

-= V02 - 4ac 21, 2 = 20

-(-4) + X1,2 = (-4)2 – 4. 1. 13 2.1

4+ 16 52 T1, 2=-

4 = V-36 11,2 =

44 62 T1, 2=-

-21,2 = 2 +31

.

total 3 eigenvalues are

I=0, I = 2+3i, I = 2 – 31

.

.

for r=0 (A- rl) is,

s -1 9 \ 100) -9 6 -15 -0.10 10 {-6 4 -10/ to 01/ -1

78 -9 -6 -1 6 4 9 \ -15 -10/

R1 + R2

7-9 18 (-6 6 -1 4 -15} 9 -10/

R2 + R2+.. R

R3 + R3- R1

7-9 6 = 10 1 [ 0 0 -15) 13 0 /

R2 + R2

7-9 6 | 0 1 (0 0 -15) -1 0

R1 + R1 - 6. Ry

7-9 | 0 0 0 1 0 -9 -1 0/

R1 + RI

10 1 [01-1 (0 0 0/

reduced system is

10 1 [01-11 [0 0 0/

sz=fi 12- = 1 .. JO=2-RU 10=2+

2= 2.................... free

.

general solution is

11

take z=1

\begin{pmatrix}x\\ y\\ z\end{pmatrix}=\begin{pmatrix}-1\\ 1\\ 1\end{pmatrix}

eigenvector is

V1 = 1 dimension : 1

..

.

for I = 2 +31 (A- 1) is.

18 -1 -9 6 (-6 4 9 (100) -15 | - (2 + 3i) 1 0 1 0 -10/ 1001)

(6 – 3i - 1 - 94-3i -6 4 9 -15 -12 - 31/

R$ * *3-(3-3): R2 + R2 - 1

R3 R3 -

9 9 76 - 3i 0 -1 14 - i18 | 127 = - 21 + +

R2 HR

=\begin{pmatrix}6-3i&-1&9\\ 0&\frac{16}{5}-i\frac{2}{5}&-\frac{24}{5}+i\frac{3}{5}\\ 0&\frac{14}{5}-i\frac{18}{5}&-\frac{21}{5}+i\frac{27}{5}\end{pmatrix}

R_3\:\leftarrow \:R_3-\left(1-i\right)\cdot \:R_2

/6 — 3i = | о \ o -1 16 — 2 o 9 | — 4 +і! o |

\:R_2\:\leftarrow \left(\frac{4}{13}+i\frac{1}{26}\right)\cdot \:R_2

=\begin{pmatrix}6-3i&-1&9\\ 0&1&-\frac{3}{2}\\ 0&0&0\end{pmatrix}

R_1\:\leftarrow \:R_1+1\cdot \:R_2

=\begin{pmatrix}6-3i&0&\frac{15}{2}\\ 0&1&-\frac{3}{2}\\ 0&0&0\end{pmatrix}

R_1\:\leftarrow \left(\frac{2}{15}+i\frac{1}{15}\right)\cdot \:R_1

=\begin{pmatrix}1&0&1+i\frac{1}{2}\\ 0&1&-\frac{3}{2}\\ 0&0&0\end{pmatrix}

reduced system is

\begin{pmatrix}1&0&1+i\frac{1}{2}\\ 0&1&-\frac{3}{2}\\ 0&0&0\end{pmatrix}\begin{pmatrix}x\\ y\\ z\end{pmatrix}=\begin{pmatrix}0\\ 0\\ 0\end{pmatrix}

\begin{Bmatrix}x+\left(1+i\frac{1}{2}\right)z=0\\ y-\frac{3}{2}z=0\end{Bmatrix}.....................\begin{Bmatrix}x=-\left(1+i\frac{1}{2}\right)z\\ y=\frac{3}{2}z\end{Bmatrix}

2= 2.................... free

.

general solution is,

\begin{pmatrix}x\\ y\\ z\end{pmatrix}=\begin{pmatrix}-\left(1+i\frac{1}{2}\right)z\\ \frac{3}{2}z\\ z\end{pmatrix}

take z=2

– 2 أ \ /

eigenvector is

dimension : 1

.

.

now to find the eigenspace for x=2-3i,

just change the sign of i (complex part)

so third eigenspace is

-2+i 3 V3 = dimension : 1

.

.

.

final answer is

for dy = 0, 0,- () dimension for 11 = 0, 01 = dimension : 1

for \:\:\:\:\lambda_2=2+3i, \:\:\:\:\:\:{\color{Red} v_2=\begin{pmatrix}-2-i\\ 3\\ 2\end{pmatrix}}\:\:\:\:\:dimension:\:1

for \:\:\:\:\lambda_3=2-3i, \:\:\:\:\:\:{\color{Red} v_3=\begin{pmatrix}-2+i\\ 3\\ 2\end{pmatrix}}\:\:\:\:\:dimension:\:1